A man traveled at a speed of 20km/h for 30 minutes. How far did he travel?
Answer by nerdybill(6962)

A man traveled at a speed of 20km/h for 30 minutes. How far did he travel?
convert 30 minutes to hours:
30(1/60) = 30/60 = 1/2
.
apply distance formula: d=rt
20(1/2)
20/2
10 km